# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 5.2. Twin graphs and trees of thick $\theta$-graphs
We now prove the final part of the lemma. Let $u$ be another essential vertex of $K_{F_{0}}^{\ast}$. As above, consider finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$ and decompose $K_{\hat{F}}^{\ast}$ as $K_{F}^{\ast}\#K_{t}^{\ast}$. Since $\theta$-graphs are $2$-connected, if the connected sum operation is performed away from $u_{F}$ and $v_{F}$ then each component of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ has connected preimage in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$. Moreover, the inclusion in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$ of the preimage of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is bijective on components. If the connected sum operation is performed at $u_{F}$ or $v_{F}$, say at $v_{F}$, then $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is a subspace of $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$. The closure of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ where $P_{v_{F}}$ is the blow-up divisor of $v_{F}$. We see that $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is obtained from $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ by gluing a copy of the half-open interval $[0,1)$ along 00 to each point in $P_{v_{F}}$. Thus $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ has a deformation retraction onto $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ and it follows that the inclusion of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}$ is bijective on components. Let $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ be a cofinal sequence of finite b-subtrees and, for each $n$, let $X_{n}=K_{F_{n}}^{\ast}\setminus\{u_{F_{n}},v_{F_{n}}\}$ and let $\tilde{X}_{n}$ be the preimage in $Z$ of $X_{n}$. Then $Z\setminus\{\tilde{u},\tilde{v}\}$ is the ascending union of the $\tilde{X}_{n}$. By Lemma 4.13, the preimage in $Z$ of any component of any $X_{n}$ is connected. Consequently, the preimage in $Z$ of any component of $X_{0}=K_{F_{0}}^{\ast}\setminus\{u,v\}$ is contained in a component of $Z\setminus\{\tilde{u},\tilde{v}\}$. By the above argument, the inclusion of $\tilde{X}_{n}$ in $\tilde{X}_{n+1}$ induces a bijection on components. It follows that the map $\pi_{0}(K_{F_{0}}^{\ast}\setminus\{u,v\})\to\pi_{0}(Z\setminus\{\tilde{u},\tilde{v}\})$ sending a component $C$ of $K_{F_{0}}^{\ast}\setminus\{u,v\}$ to the component of $Z\setminus\{\tilde{u},\tilde{v}\}$ that contains the preimage of $C$ is bijective. ∎
@@
Recall the description of Bowditch classes in Section 3.1: a pair of points $\{x,y\}$ is a Bowditch cut pair if $\deg(x)=\deg(y)\geq 2$ and the common degree of $x$ and $y$ is equal to the number of connected components in the complement of $\{x,y\}$. The degree of $\{x,y\}$ is defined as the common degree of $x$ and $y$. In the case of graphs we have used the terminolgoy twins pairs for Bowditch cut pairs. We will also consider Bowditch cut pairs in the case of trees of punctured thick $\theta$-graphs.
@@
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$. Let $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ be disjoint Bowditch cut pairs of $Z$ such that $\{x,y\}$ is not a Bowditch cut pair of $Z$. If $\{x,x^{\prime}\}$ has degree at least $3$ and $\{y,y^{\prime}\}$ has degree at least $3$ then there exists a Bowditch cut pair $\{z,z^{\prime}\}$ of degree $2$ such that $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ are contained in distinct components of $Z\setminus\{z,z^{\prime}\}$.
@@
By Corollary 5.32, there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projection $\{x_{F},x^{\prime}_{F}\}$ of $\{x,x^{\prime}\}$ in the reduced partial union $K_{F}^{\ast}$ is a twin pair of the same degree as $\{x,x^{\prime}\}$ and the projection $\{y_{F},y^{\prime}_{F}\}$ of $\{y,y^{\prime}\}$ in $K_{F}^{\ast}$ is a twin pair of the same degree as $\{y,y^{\prime}\}$. Then $x,x^{\prime},y,y^{\prime}$ are the canonical lifts of $x_{F},x^{\prime}_{F},y_{F},y^{\prime}_{F}$ given by Lemma 5.30 and so $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ must be distinct twin pairs of $K_{F}^{\ast}$. Recall that $K_{F}^{\ast}$ is obtained from thick $\theta$-graphs by a series of connected sum operations. By first performing the connected sum operations at essential vertices, we can assume that the connected sum operations take place only at degree $2$ points so that, by Proposition 5.19(1), the connected sum operations preserve essential twins. Then $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ correspond to twin pairs in distinct thick $\theta$-graph summands and, consequently, some blow-up divisor $\{q_{1},q_{2}\}=P_{q}\subset K_{F}^{\ast}$ of a degree $2$ point $q$ at which a connected sum operation was performed separates $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ into distinct components. Since connected sums of thick $\theta$-graphs are 2-connected, the complement $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$ has two components. Since $q_{1}$ and $q_{2}$ are stable points (i.e. no connected sum operation is performed at these points to obtain a larger reduced partial union) they have singleton preimages $\{\tilde{q}_{1}\}$ and $\{\tilde{q}_{2}\}$ in $Z$ and $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ is the preimage in $Z$ of $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$. Then Lemma 4.13 implies that $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ has $2$ components and Lemma 5.26 implies that $\tilde{q}_{1}$ and $\tilde{q}_{2}$ have degree $2$ so that we can take $\{z,z^{\prime}\}=\{\tilde{q}_{1},\tilde{q}_{2}\}$. ∎
@@
[…]
@@
Let $T$ be the underlying tree of the tree system of graphs whose limit is homeomorphic to $\partial G$. We identify $\partial G$ with this limit throughout the proof.
@@
Take $x\in\zeta$ an $y\in\eta$. By Lemma 5.28, there are at most two components of $\partial G\setminus\{x,y\}$. But $\zeta$ and $\eta$ are distinct Bowditch classes so the points $x$ and $y$ do not form a Bowditch cut pair. Thus $\partial G\setminus\{x,y\}$ does not have two components and so it must be connected. By Lemma 5.29, there is a finite b-subtree $F\subset T$ for which the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ in the reduced partial union $K_{F}^{\ast}$ have degree $2$ and the complement $K_{F}^{\ast}\setminus\{x_{F},y_{F}\}$ is connected. By Theorem 5.24, the reduced partial sum $K_{F}^{\ast}$ is a twin graph. By Lemma 5.25, there is an essential twin pair $v,\bar{v}$ such that $x_{F}$ and $y_{F}$ are contained in distinct components of $K_{F}^{\ast}\setminus\{v,\bar{v}\}$. Then, by Corollary 5.31, the pairs $x$ and $y$ are contained in distinct components $C_{x}$ and $C_{y}$ of the complement $\partial G\setminus\{z,z^{\prime}\}$ of a Bowditch cut pair $\{z,z^{\prime}\}$ in $\partial G$ of degree $\deg(v)\geq 3$. Then, by Lemma 5.22, we have $\zeta\subset C_{x}$ and $\eta\subset C_{y}$, contradicting the fact that $\zeta$ and $\eta$ are not separated. ∎
@@ 6. Regular trees of graphs
In this section we describe a class of tree systems of punctured graphs which depend uniquely on certain finite collections of defining data (which we call graphical connecting systems, see Definition 6.1 below). We call the limits of such tree systems the regular trees of graphs. Although we do not justify this fully in the present paper, the regular trees of graphs belong to a larger class of spaces called Markov compacta, described by certain recursive procedures based on some finite collections of data (see [11, 18]). Since, due to the main result from [18], Gromov boundaries of hyperbolic groups are all homeomorphic to Markov compacta, the regular trees of graphs are good candidates for spaces that can be potentially realized as Gromov boundaries of some hyperbolic groups. Actually, the main goal of the present paper is to characterize those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are (homeomorphic to some) trees of graphs, and to show that these trees of graphs are then regular.
Recall that in this paper we allow graphs to have multiple edges and loop edges. Given a graph $\Gamma$ and its vertex $v$, the link of $\Gamma$ at $v$, denoted $\operatorname{Lk}(v)$, is the set of oriented edges of $\Gamma$ issuing from $v$. Equivalently, $\operatorname{Lk}(v)$ can be described as the blow-up divisor $P_{v}$ of the blow-up $|\Gamma|^{\#}(v)$ at $v$ (as described in Subsection 4.1).
Given an oriented edge $\varepsilon$ of a graph $\Gamma$, we denote by $\bar{\varepsilon}$ the same edge taken with the opposite orientation.
@@
    $\Gamma=\sqcup_{j=1}^{k}\Gamma_{j}$ is the disjoint union of a finite nonempty collection of nontrivial connected finite graphs $\Gamma_{j}$ (in particular, $k\geq 1$);
@@
    $T$ is the unique tree bipartitioned into black and white vertices such that the order of any white vertex is 2, and the order of any black vertex is countable infinite;
@@
      if the image $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is a V-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, and if it corresponds to a vertex $v\in V_{\Gamma^{\circ}_{j(t_{1})}}$, then $\Gamma_{j(t_{2})}$ is this component of $\Gamma$ which contains the vertex $a(v)$, and the image $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is the V-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$ corresponding to $a(v)$; moreover, the composition map $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$ coincides with the bijection $\alpha_{v}$ (after natural identifications of the vertex links with the sets of elements of the V-peripherals, at the vertices $v$ and $a(v)$);
@@
      if $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, then $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$; moreover, if $e_{i}$ is the edge of $\Gamma$ containing $\varphi_{[u,t_{i}]}(\Sigma_{u})$, for $i=1,2$, and if $\varepsilon_{1}$ is any oriented edge with $|\varepsilon_{1}|=e_{1}$, then denoting by $\varepsilon_{2}$ this oriented edge related to $e_{2}$ whose orientation is induced from that of $\varepsilon_{1}$ via $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$, we have $(\varepsilon_{1},\varepsilon_{2})\in\mathcal{A}$;
@@
    for any black vertex $t$ of $T$ we have
    1. (a)
      for each $v\in V_{\Gamma_{j(t)}}$ there is $u\in N_{t}$ such that $\varphi_{[u,t]}(\Sigma_{u})$ is the V-peripheral corresponding to $v$, and
    2. (b)
      for any oriented edge $\varepsilon$ of $\Gamma_{j(t)}$, denoting by $\Omega(\varepsilon)$ the family of those oriented E-peripherals in $K_{t}\cong\Gamma^{\circ}_{j(t)}$ which are contained in $|\varepsilon|$, and which are oriented consistently with the orientation of $\varepsilon$, and putting $A_{\varepsilon}:=\{\varepsilon^{\prime}:(\varepsilon,\varepsilon^{\prime})\in\mathcal{A}\}$, consider the map $f_{\varepsilon}:\Omega(\varepsilon)\to A_{\varepsilon}$ which to any E-peripheral $\omega\in\Omega(\varepsilon)$, $\omega=\varphi_{[u,t]}(\Sigma_{u})$, associates this oriented edge $\varepsilon^{\prime}$ of $\Gamma$ for which $\varphi_{[u,t^{\prime}]}(\Sigma_{u})\subset|\varepsilon^{\prime}|$ (where $t^{\prime}$ is the black vertex of $T$ adjacent to $u$ and distinct from $t$) and which is oriented in the way induced by the map $\varphi_{[u,t^{\prime}]}\circ\varphi_{[u,t]}^{-1}$ from the orientation of $\varepsilon$; then for each $\varepsilon^{\prime}\in A_{\varepsilon}$ the union $\bigcup f_{\varepsilon}^{-1}(\varepsilon^{\prime})$ is dense (and in particular nonempty) in $|\varepsilon|\cap\Gamma^{\circ}$.
Though technically involved, the above definitions are in fact quite natural, and their significance stems from the following lemma, in view of which to any graphical connecting system $\mathcal{R}$ there is associated some unique up to isomorphism tree system of punctured graphs (which we denote $\Theta[{\mathcal{R}}]$), and some unique up to homeomorphism topological space having the form of a tree of graphs (which we denote $\mathcal{X}(\mathcal{R})$).

### v2

@@ 5.3. Proof of the impliciation (1$\Rightarrow$(3) of Theorem B
We now prove the final part of the lemma. Let $u$ be another essential vertex of $K_{F_{0}}^{\ast}$. As above, consider finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$ and decompose $K_{\hat{F}}^{\ast}$ as $K_{F}^{\ast}\#K_{t}^{\ast}$. Since $\theta$-graphs are $2$-connected, if the connected sum operation is performed away from $u_{F}$ and $v_{F}$ then each component of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ has connected preimage in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$. Moreover, the inclusion in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$ of the preimage of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is bijective on components. If the connected sum operation is performed at $u_{F}$ or $v_{F}$, say at $v_{F}$, then $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is a subspace of $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$. The closure of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ where $P_{v_{F}}$ is the blow-up divisor of $v_{F}$. We see that $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is obtained from $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ by gluing a copy of the half-open interval $[0,1)$ along 00 to each point in $P_{v_{F}}$. Thus $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ has a deformation retraction onto $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ and it follows that the inclusion of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}$ is bijective on components. Let $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ be a cofinal sequence of finite b-subtrees such that for each $n\geq 0$ no b-subtree lies properly between $F_{n}$ and $F_{n+1}$. For each $n$, let $X_{n}=K_{F_{n}}^{\ast}\setminus\{u_{F_{n}},v_{F_{n}}\}$ and let $\tilde{X}_{n}$ be the preimage in $Z$ of $X_{n}$. Then $Z\setminus\{\tilde{u},\tilde{v}\}$ is the ascending union of the $\tilde{X}_{n}$. By Lemma 4.13, the preimage in $Z$ of any component of any $X_{n}$ is connected. Consequently, the preimage in $Z$ of any component of $X_{0}=K_{F_{0}}^{\ast}\setminus\{u,v\}$ is contained in a component of $Z\setminus\{\tilde{u},\tilde{v}\}$. By the above argument, the inclusion of $\tilde{X}_{n}$ in $\tilde{X}_{n+1}$ induces a bijection on components. It follows that the map $\pi_{0}(K_{F_{0}}^{\ast}\setminus\{u,v\})\to\pi_{0}(Z\setminus\{\tilde{u},\tilde{v}\})$ sending a component $C$ of $K_{F_{0}}^{\ast}\setminus\{u,v\}$ to the component of $Z\setminus\{\tilde{u},\tilde{v}\}$ that contains the preimage of $C$ is bijective. ∎
@@
Recall the description of Bowditch classes in Section 3.1: a pair of points $\{x,y\}$ is a Bowditch cut pair if $\deg(x)=\deg(y)\geq 2$ and the common degree of $x$ and $y$ is equal to the number of connected components in the complement of $\{x,y\}$. The degree of $\{x,y\}$ is defined as the common degree of $x$ and $y$. In the case of graphs we have used the terminology twin pairs for Bowditch cut pairs. We will also consider Bowditch cut pairs in the case of trees of punctured thick $\theta$-graphs.
@@
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$, and let $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ be disjoint Bowditch cut pairs of $Z$. If $\{x,x^{\prime}\}$ has degree at least $3$ and $\{y,y^{\prime}\}$ has degree at least $3$ then there exists a Bowditch cut pair $\{z,z^{\prime}\}$ of degree $2$ such that $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ are contained in distinct components of $Z\setminus\{z,z^{\prime}\}$.
@@
It follows from Corollary 5.32 that there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projection $\{x_{F},x^{\prime}_{F}\}$ of $\{x,x^{\prime}\}$ in the reduced partial union $K_{F}^{\ast}$ is a twin pair of the same degree as $\{x,x^{\prime}\}$ and the projection $\{y_{F},y^{\prime}_{F}\}$ of $\{y,y^{\prime}\}$ in $K_{F}^{\ast}$ is a twin pair of the same degree as $\{y,y^{\prime}\}$. Then $x,x^{\prime},y,y^{\prime}$ are the canonical lifts of $x_{F},x^{\prime}_{F},y_{F},y^{\prime}_{F}$ given by Lemma 5.30 and so $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ must be distinct essential twin pairs of $K_{F}^{\ast}$. Recall that $K_{F}^{\ast}$ is obtained from thick $\theta$-graphs by a series of connected sum operations. By first performing the connected sum operations at essential vertices, we can assume that the connected sum operations take place only at degree $2$ points so that, by Proposition 5.19(1), the connected sum operations preserve essential twins. Then $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ correspond to essential twin pairs in distinct thick $\theta$-graph summands and, consequently, some blow-up divisor $\{q_{1},q_{2}\}=P_{q}\subset K_{F}^{\ast}$ of a degree $2$ point $q$ at which a connected sum operation was performed separates $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ into distinct components. Since connected sums of thick $\theta$-graphs are 2-connected, the complement $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$ has two components. Since $q_{1}$ and $q_{2}$ are stable points (i.e. no connected sum operation is performed at these points to obtain a larger reduced partial union) they have singleton preimages $\{\tilde{q}_{1}\}$ and $\{\tilde{q}_{2}\}$ in $Z$ and $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ is the preimage in $Z$ of $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$. Then Lemma 4.13 implies that $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ has $2$ components and Lemma 5.26 implies that $\tilde{q}_{1}$ and $\tilde{q}_{2}$ have degree $2$ so that we can take $\{z,z^{\prime}\}=\{\tilde{q}_{1},\tilde{q}_{2}\}$. ∎
@@
We are now ready to prove the main result of this section: the implication (1$\Rightarrow$(3) of Theorem B. We restate and prove the implication below.
@@
Let $T$ be the underlying tree of the tree system of punctured graphs whose limit is homeomorphic to $\partial G$. We identify $\partial G$ with this limit throughout the proof.
@@
Take $x\in\zeta$ an $y\in\eta$. By Lemma 5.28, there are at most two components of $\partial G\setminus\{x,y\}$. But $\zeta$ and $\eta$ are distinct Bowditch classes so the points $x$ and $y$ do not form a Bowditch cut pair. Thus $\partial G\setminus\{x,y\}$ does not have two components and so it must be connected. By Lemma 5.29, there is a finite b-subtree $F\subset T$ for which the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ in the reduced partial union $K_{F}^{\ast}$ have degree $2$ and the complement $K_{F}^{\ast}\setminus\{x_{F},y_{F}\}$ is connected. By Theorem 5.24, the reduced partial sum $K_{F}^{\ast}$ is a twin graph. By Lemma 5.25, there is an essential twin pair $v,\bar{v}$ such that $x_{F}$ and $y_{F}$ are contained in distinct components of $K_{F}^{\ast}\setminus\{v,\bar{v}\}$. Then, by Corollary 5.31, the points $x$ and $y$ are contained in distinct components $C_{x}$ and $C_{y}$ of the complement $\partial G\setminus\{z,z^{\prime}\}$ of a Bowditch cut pair $\{z,z^{\prime}\}$ in $\partial G$ of degree $\deg(v)\geq 3$. Since moreover, by Lemma 5.22, we have $\zeta\subset C_{x}$ and $\eta\subset C_{y}$, this yields a contradiction with the fact that $\zeta$ and $\eta$ are not separated. ∎
@@ 6. Regular trees of graphs
In this section we describe a class of tree systems of punctured graphs which depend uniquely on certain finite collections of defining data (which we call graphical connecting systems, see Definition 6.1 below). We call the limits of such tree systems the regular trees of graphs. Although we do not justify this in the present paper (since this is not essential for our results), the regular trees of graphs belong to a larger class of spaces called Markov compacta, which are described by certain recursive procedures based on some finite collections of data (see [Dra07, Paw15, Cza22]). Since, due to the main result from [Paw15], Gromov boundaries of hyperbolic groups are all homeomorphic to Markov compacta, the regular trees of graphs are good candidates for spaces that can be potentially realized as Gromov boundaries of some hyperbolic groups. Actually, the main goal of the present paper is to characterize those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are (homeomorphic to some) trees of graphs, and to show that these trees of graphs are then regular.
Recall that in this paper we allow graphs to have multiple edges and loop edges. Given a graph $\Gamma$ and its vertex $v$, the link of $\Gamma$ at $v$, denoted $\operatorname{Lk}(v)$, is the set of oriented edges of $\Gamma$ issuing from $v$. Equivalently, $\operatorname{Lk}(v)$ can be described as the blow-up divisor $P_{v}$ of the blow-up $|\Gamma|^{\#}(v)$ at $v$ (as described in Section 4.1).
Given an oriented edge $\varepsilon$ of a graph $\Gamma$, we denote by $\bar{\varepsilon}$ the same edge taken with the opposite orientation; we denote by $|\varepsilon|$ the unoriented edge which underlies the oriented edge $\varepsilon$.
@@
    $\Gamma=\sqcup_{j=1}^{k}\Gamma_{j}$ is the disjoint union of a finite nonempty collection of connected finite graphs $\Gamma_{j}$, each of which is nonempty and not equal to a single vertex (in particular, $k\geq 1$);
@@
Before presenting the next crucial definition (Definition 6.2) we need some further terminological and notational preparations. Let ${\mathcal{N}}_{\Gamma}$ be a standard family of normal neighbourhoods of $\Gamma$, as described in Example 2.2, and let $\Gamma^{\circ}:=|\Gamma|\setminus\bigcup\{\operatorname{int}B:B\in{\mathcal{N}}_{\Gamma}\}$ be the space with peripherals, called the punctured graph $\Gamma$, as described in the same example. Denote by $\Omega^{E}_{\Gamma}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}=\emptyset\}$ the set of those peripherals in $\Gamma^{\circ}$ which correspond to those $B\in{\mathcal{N}}_{\Gamma}$ which are normal neighbourhoods of interior points of the edges of $\Gamma$. Similarly, put $\Omega_{\Gamma}^{V}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}\neq\emptyset\}$. We will call the elements of the above sets E-peripherals and V-peripherals, respectively. Observe that each E-peripheral is a set of cardinality 2, i.e., a doubleton. Each such set can be equipped with one of the two possible orders, so that it becomes an ordered pair of points, and we will call any such order an orientation associated to this E-peripheral. We denote by $\widetilde{\Omega}^{E}_{\Gamma}$ the set of all oriented E-peripherals of $\Gamma^{\circ}$, i.e., the set $\{(p,q):\{p,q\}\in\Omega^{E}_{\Gamma}\}$. Observe that each of the oriented E-peripherals may be either compatible or incompatible with the fixed orientation of the edge $e$ of $\Gamma$ in the interior of which it is contained. Actually, in the definition below we use the convention of identifying orientations of E-peripherals of $\Gamma^{\circ}$ with the corresponding compatible orientations of the edges of $\Gamma$ containing those peripherals. For any oriented peripheral $\delta\in\widetilde{\Omega}^{E}_{\Gamma}$, we denote by $|\delta|$ the corresponding underlying (unoriented) peripheral, and by $\bar{\delta}$ the same peripheral $|\delta|$ oriented oppositely to $\delta$.
@@
    $T$ is the unique tree bipartitioned into black and white vertices such that the degree of any white vertex is 2, and the degree of any black vertex is countable infinite;
@@
      if the image $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is a V-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, and if it corresponds to a vertex $v\in V_{\Gamma_{j(t_{1})}}$, then $\Gamma_{j(t_{2})}$ is the component of $\Gamma$ which contains the vertex $a(v)$, and the image $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is the V-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$ corresponding to $a(v)$; moreover, the composition map $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$ coincides with the bijection $\alpha_{v}$ (after natural identifications of the vertex links with the sets of elements of the V-peripherals, at the vertices $v$ and $a(v)$);
@@
      if $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, then $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$; moreover, if $e_{i}$ is the edge of $\Gamma$ containing $\varphi_{[u,t_{i}]}(\Sigma_{u})$, for $i=1,2$, and if $\varepsilon_{1}$ is any oriented edge with $|\varepsilon_{1}|=e_{1}$, then denoting by $\varepsilon_{2}$ the oriented edge related to $e_{2}$ whose orientation is induced from that of $\varepsilon_{1}$ via $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$, we have $(\varepsilon_{1},\varepsilon_{2})\in\mathcal{A}$;
@@
    for any black vertex $t$ of $T$ and for any oriented edge $\varepsilon$ of $\Gamma_{j(t)}$, denote by $\widetilde{\Omega}(\varepsilon)$ the family of those oriented E-peripherals in $K_{t}\cong\Gamma^{\circ}_{j(t)}$ which are contained in $|\varepsilon|$, and which are oriented consistently with the orientation of $\varepsilon$; put also $O_{\varepsilon}:=\{\varepsilon^{\prime}:(\varepsilon,\varepsilon^{\prime})\in\mathcal{A}\}$, and consider the map $f_{\varepsilon}:\widetilde{\Omega}(\varepsilon)\to O_{\varepsilon}$ described as follows: if $\omega\in\widetilde{\Omega}(\varepsilon)$ and $u$ is the white vertex of $T$ for which $|\omega|=\varphi_{[u,t]}(\Sigma_{u})$ and $t^{\prime}$ is the black vertex of $T$ adjacent to $u$ and distinct from $t$ then $f_{\varepsilon}(\omega)$ is the oriented edge $\varepsilon^{\prime}$ of $\Gamma_{j(t^{\prime})}$ containing $\varphi_{[u,t^{\prime}]}(\Sigma_{u})$ and oriented in the way induced by the map $\varphi_{[u,t^{\prime}]}\circ\varphi_{[u,t]}^{-1}$ from the orientation of $\varepsilon$; it is required that for each $\varepsilon^{\prime}\in O_{\varepsilon}$ the union $\bigcup\{|\omega|:\omega\in\widetilde{\Omega}(\varepsilon)\hbox{ and }f_{\varepsilon}(\omega)=\varepsilon^{\prime}\}$ is dense (and in particular nonempty) in $|\varepsilon|\cap\Gamma^{\circ}$.
Though technically involved, the above definitions (i.e. Definitions 6.1 and 6.2) are in fact quite natural, and their significance stems from the following lemma, in view of which to any graphical connecting system $\mathcal{R}$ there is associated some unique up to isomorphism tree system of punctured graphs (which we denote $\Theta[{\mathcal{R}}]$), and some unique up to homeomorphism topological space having the form of a tree of graphs (which we denote $\mathcal{X}(\mathcal{R})$, and which is given as $\mathcal{X}(\mathcal{R}):=\lim\Theta[{\mathcal{R}}]$).